3X + 2Y= 8 X=3Y +10 HOW WOULD WE START THIS PROBLEM
Answer by Abbey(339)

Use substitution:
Since we know what x=, we can use this information to rewrite the following:
3(X) + 2Y = 8
3(3y+10) + 2Y = 8
clear the parentheses:
9y+30+2y=8
Consolidate like terms
11y+30=8
Subtract 30 from both sides:
11y=-22
Divide both sides by 11
y=-2
Substitute y back into either of your equations to find x:
3x+2(-2)=8
3x+-4=8
3x=12
x=4
Check with your other equation:
4=3(-2)+10
4=-6+10
4=4
True statement, so the solution set is (4,-2)